# Bang–Bang Optimal Control of Vaccination in Metapopulation Epidemics with Linear Cost Structures

### v2

@@ 1.  Introduction
The remainder of the paper is organized as follows. Section 2 formulates the metapopulation epidemic vaccination model as an optimal control problem, and Section 3 presents necessary optimality conditions for it. Section 4 is dedicated to the main results, namely the absence of singular arcs, the bang–bang structure, and other technical lemmas. Section 5 presents numerical experiments confirming our theoretical results, and Section 6 concludes with possible extensions.
Notation. We write $\mathbb{R}_{\geq 0}=[0,\infty)$ and $\|\cdot\|_{\infty}$ for the essential‐supremum norm. The space $L^{\infty}([0,T];\mathbb{R}^{k}_{\geq 0})$ consists of (equivalence classes of) essentially bounded measurable maps $u:[0,T]\to\mathbb{R}^{k}_{\geq 0}$. If $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$, then $D_{x}g$ or $\bigl{[}\partial g_{j}/\partial x_{i}\bigr{]}_{j,i}$ denotes its $m\times n$ Jacobian; for any matrix $M$, $M_{ij}$ is its entry in row $i$, column $j$. Given a set $S$, $\bar{S}$ is its closure, $\inf S$ its infimum, and $\sup S$ its supremum. Finally, “a.e.” abbreviates “almost everywhere.”
2.  General Epidemic Model Formulation
@@
Set $\boldsymbol{s}=(s^{(1)},\dots,s^{(K)})$, $\boldsymbol{x}=(x^{(1)}|\dots|x^{(K)})\in\mathbb{R}^{d\times K}$ and the total vaccinated population $V(t)\coloneqq\sum_{i=1}^{K}n_{i}V^{(i)}(t)$. The epidemic dynamics, subject to vaccination rates $u_{i}\colon[0,T]\to\mathbb{R}_{\geq 0}$, is governed by
$\frac{ds^{(i)}}{dt}$ $=-s^{(i)}f_{i}(\boldsymbol{x})-u_{i}s^{(i)},$ (1)
$\frac{dx^{(i)}}{dt}$ $=s^{(i)}f_{i}(\boldsymbol{x})e_{1}+g(x^{(i)}),$
$\frac{dV}{dt}$ $=\sum_{i=1}^{K}u_{i}n_{i}s^{(i)},$
@@
where $g_{j}$ refers to the $j$th component of $g$. Additionally, $g(0)=0$ guarantees that the disease-free equilibrium is invariant. For positivity, we assume $f_{i}(\boldsymbol{x})\geq 0$ if $\boldsymbol{x}\geq 0$ and $g_{i}$ is essentially non-negative, this is, if $x\in\mathbb{R}^{d},$ $x\geq 0$ and $x_{j}=0$, then $g_{j}(x)\geq 0$.
Proposition 1.
@@
where $c_{v}>0$ and $f_{0}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$. Here, $f_{0}$ maps the disease‐stage vector to a per‐individual health cost.
@@
We look for $u$ to minimize the finite‐horizon cost $J[u]$ on $[0,T]$ with no terminal constraint, so our focus is on the open-loop solution of the optimization problem (7) rather than on the closed-loop optimizer or stabilizer. Although the uncontrolled system admits a disease‐free equilibrium (and may admit endemic equilibria depending on $g$), their presence does not affect the optimization.
Proposition 2.
@@
Proof.
We show that the assumptions of the existence results from Cesari [27, Theorem 1] are satisfied for our optimal control problem (7). By Proposition 1 and condition (2), every solution of the epidemic system (1) remains in a positively invariant and bounded subset $\Omega\subset\mathbb{R}^{K(d+1)+1}$. Furthermore, the constraint $V(t)\leq D_{\varepsilon}(t)$ ensures that $\Omega$ is closed, hence compact. The cost functional (6) is continuous in both state and control variables, and the system’s right‐hand side in (1) is locally Lipschitz in $x$ and continuous in $(x,u)$. For each $t\in[0,T],x\in\Omega$, the admissible control set is defined as
$U(t,x)\coloneqq\{\,u\in L^{\infty}\bigl{(}[0,T];\mathbb{R}^{K}_{\geq 0}\bigr{)}:0\leq u_{i}\,s^{(i)}(t)\leq v_{i}^{\max}\},$
which is compact (by the continuity of the mixed control-state constraints and equation (4)) and upper semi-continuous as a set-valued map. Finally, define
$\tilde{Q}(t,x)\coloneqq\{(z^{0},z):z^{0}\geq\ell(x,u),\;z=F(x,u),\;u\in U(t,x)\},$
where $\ell$ and $F$ denote the running cost and the dynamics, respectively. Since $\ell$ and $F$ are affine and continuous in $u$ and $U(t,x)$ is convex, each $\tilde{Q}(t,x)$ is convex in $(z^{0},z)$.
We conclude that the existence of a globally optimal process is guaranteed by [27, Theorem 1]. ∎
3.  Necessary Conditions for Optimality
@@
$H$ $\coloneqq-\sum_{i=1}^{K}\psi^{s}_{i}\left(s^{(i)}\left(f_{i}(\boldsymbol{x})+u_{i}\right)\right)+\sum_{i=1}^{K}\langle\psi^{x^{(i)}},s^{(i)}f_{i}(\boldsymbol{x})e_{1}+g(x^{(i)})\rangle+\psi^{V}\sum_{i=1}^{K}u_{i}n_{i}s^{(i)}$ (8)
@@
$\dot{\psi}^{x^{(i)}}$ $=\sum_{j=1}^{K}(\psi_{j}^{s}-\psi_{1}^{x^{(j)}})s^{(j)}\left(\frac{\partial f_{j}(\boldsymbol{x})}{\partial x^{(i)}}\right)^{\top}-\psi^{x^{(i)}}D_{x}g(x^{(i)})+\lambda_{0}n_{i}\nabla f_{0}\left(x^{(i)}\right)^{\top},$
@@ 4.  Main Results
Intuitively, $\lvert\psi^{s}_{i}\rvert<\psi^{x^{(i)}}_{1}$ since, under the shadow‐price interpretation [12], infection incurs a higher cost than susceptibility. Lemma 1 proves this fact rigorously under the following simplifying assumptions.
Assumption 1.
@@
    The state‐dependent cost is nontrivial. In particular, if we define
    $\begin{split}I&\coloneqq\Bigl{\{}j\in\{2,\dots,d\}\;\Bigm{|}\;\exists\,m\geq 1,(i_{k})_{k=0,\dots,m}\text{ s. t. }i_{0}=i,i_{m}=j,\text{ and }G_{i_{k+1},i_{k}}>0,\text{ for all }k\Bigr{\}}.\end{split}$
@@
Proof.
See the Appendix (page APPENDIX). ∎
@@
Theorem 2 (Optimal vaccination structure).
@@
5.  Numerical experiments
We validate our theoretical findings by simulating optimal vaccination strategies in three cases (networks with $3$, $5$, and $8$ cities, each city following a simple SIR model with daily commuting between them). All simulations share a 28‑day horizon ($T=28$), recovery rate $\gamma=1/7$ day-1, and home‑stay fraction $\alpha=0.64$ (workers spend $1-\alpha$ of each week in the destination city)..
@@ 5.  Numerical experiments
We also solved the problem, but with a quadratic vaccination cost. Figure 2 indicates that the quadratic objective yields a smoother control that approximates the bang–bang profile; the resulting cost difference is negligible (bottom-right panel). The control resulting from the quadratic cost is harder to implement in practice than the bang-bang given in the linear-cost case.
@@ 6.  Conclusion
Proof of Lemma 1.
Define the variables $z^{(i)}(t)\in\mathbb{R}$, $y^{(i)}(t)\in\mathbb{R}^{d}$, $w^{(i)}(t)\in\mathbb{R}^{d}$, and $\varphi^{(i)}(t)\in\mathbb{R}^{d}$ as:
$z^{(i)}(t)$ $\coloneqq\psi_{i}^{s}(T-t)-\psi_{1}^{x^{(i)}}(T-t),$
$y^{(i)}(t)$ $\coloneqq-\psi^{x^{(i)}}(T-t)^{\top},$
$w^{(i)}(t)$ $\coloneqq\sum_{j=1}^{K}z^{(j)}(t)s^{(j)}(T-t)\beta_{ij}+cn_{i},$
$\varphi^{(i)}(t)$ $\coloneqq w^{(i)}(t)+G^{T}y^{(i)}(t),$
and note that the initial conditions are $z^{(i)}(0)=y^{(i)}(0)=0$, $w^{(i)}(0)=cn_{i}$ and $\varphi^{(i)}(0)=cn_{i}$. Using the definitions, these variables satisfy the system
@@
ACKNOWLEDGMENT
The second author acknowledges the financial support of the Brazilian agencies FAPERJ processes E-26/203.223/2017 and E-26/201.346/2021 and CNPq process 312407/2023-8. The current research was mainly developed while the first author was a Master’s student at FGV EMAp, with institutional financial support.

### v3

@@ I Introduction
The remainder of the paper is organized as follows. Section II formulates the metapopulation epidemic vaccination model as an optimal control problem, and Section III presents necessary optimality conditions for it. Section IV is dedicated to the main results, namely the absence of singular arcs, the bang-bang structure, and other technical lemmas. Section V presents numerical experiments confirming our theoretical results, and Section VI concludes with possible extensions.
Notation. We write $\mathbb{R}_{\geq 0}=[0,\infty)$ and $\|\cdot\|_{\infty}$ for the essential-supremum norm. The space $L^{\infty}([0,T];\mathbb{R}^{k}_{\geq 0})$ consists of (equivalence classes of) essentially bounded measurable maps $u:[0,T]\to\mathbb{R}^{k}_{\geq 0}$. If $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$, then $D_{x}g$ or $\bigl{[}\partial g_{j}/\partial x_{i}\bigr{]}_{j,i}$ denotes its $m\times n$ Jacobian; for any matrix $M$, $M_{ij}$ is its entry in row $i$, column $j$. Given a set $S$, $\bar{S}$ is its closure, $\inf S$ its infimum, and $\sup S$ its supremum. Finally, “a.e.” abbreviates “almost everywhere.”
II General Epidemic Model Formulation
@@
Set $\boldsymbol{s}=(s^{(1)},\dots,s^{(K)})$, $\boldsymbol{x}=(x^{(1)}|\dots|x^{(K)})\in\mathbb{R}^{d\times K}$ and the total vaccinated population $V(t)\coloneqq\sum_{i=1}^{K}n_{i}V^{(i)}(t)$.
The epidemic dynamics, subject to vaccination rates $u_{i}\colon[0,T]\to\mathbb{R}_{\geq 0}$, is governed by
$\dot{s}^{(i)}$ $=-s^{(i)}f_{i}(\boldsymbol{x})-u_{i}s^{(i)},$ (1)
$\dot{x}^{(i)}$ $=s^{(i)}f_{i}(\boldsymbol{x})e_{1}+g(x^{(i)}),$
$\dot{V}$ $=\sum_{i=1}^{K}u_{i}n_{i}s^{(i)},$
@@
where $g_{j}$ refers to the $j$th component of $g$. Additionally, $g(0)=0$ guarantees that the disease-free equilibrium is invariant. For positivity, we assume $f_{i}(\boldsymbol{x})\geq 0$ if $\boldsymbol{x}\geq 0$ and $g_{i}$ is essentially non-negative, this is, if $x\in\mathbb{R}^{d}$, $x\geq 0$, and $x_{j}=0$, then $g_{j}(x)\geq 0$.
Remark 1
Propositions 1, 2 and 3 remain valid if the susceptible dynamics include any Lipschitz-continuous migration term $q_{i}(\mathbf{s})$ to the $s^{(i)}$-dynamics. The bang–bang proof, however, relies on monotonicity, so extending Theorems 1 and 2 to include $q_{i}$ is left for future work. Numerical tests with linear $q_{i}$ (see Figure 2) also show one switch per week, suggesting the qualitative behavior is robust.
Proposition 1
@@
where $c_{v}>0$ and $f_{0}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$. Here, $f_{0}$ maps the disease-stage vector to a per-individual health cost.
@@
We look for $u$ to minimize the finite-horizon cost $J[u]$ on $[0,T]$ with no terminal constraint, so our focus is on the open-loop solution of the optimization problem (7) rather than on the closed-loop optimizer or stabilizer. Although the uncontrolled system admits a disease-free equilibrium (and may admit endemic equilibria depending on $g$), their presence does not affect the optimization.
Proposition 2
@@
Proof:
We check directly that all the assumptions of Cesari’s existence theorem [27, Th. 1] are fulfilled by problem (7). Let $w(t)\coloneqq(\boldsymbol{s}(t),\boldsymbol{x}(t),V(t))$.
First, define the pure-state constraint $g(t,w)\coloneqq V-D_{\varepsilon}(t)$ and the set
$\Omega\;=\;\{(t,w)\in[0,T]\times\mathbb{R}_{\geq 0}^{K(d+1)+1}\;:\;g(t,w)\leq 0\}.$
Because $D_{\varepsilon}$ is continuous and the state space is bounded by the population balance, $\Omega$ is compact. By Proposition 1 every admissible trajectory of (1) with $V\leq D_{\varepsilon}$ stays in $\Omega$.
For $(t,w)\in\Omega$, set
$U(t,w)\;=\;\Bigl{\{}u\in\mathbb{R}_{\geq 0}^{K}\;:\;m(w,u)\leq 0\Bigr{\}},$
with $m(w,u)\;=\;\bigl{(}u_{i}\,s^{(i)}-v_{i}^{\max}\bigr{)}_{i=1}^{K}$. Since $s^{(i)}(t)\geq l_{i}>0$, each component of $u$ is bounded by $v_{i}^{\max}/l_{i}$; hence $U(t,w)$ is a closed, bounded and convex box, and the continuity of $m$ implies upper semi-continuity of the map $(t,w)\mapsto U(t,w)$.
The $C^{2}$-regularity of $f_{i}$ and $g$ makes $F$ a Carathéodory mapping, locally Lipschitz in $w$ and continuous in $(w,u)$. The running cost is $\ell(w)=\sum_{i=1}^{K}n_{i}f_{0}(x^{(i)})$, and the Bolza terminal term $c_{v}V(T)$ is converted into the Lagrange form $\psi(w,u)=c_{v}\sum_{i=1}^{K}n_{i}u_{i}s^{(i)}$. Both $\ell$ and $\psi$ are continuous and affine in $u$.
For Cesari’s convexity requirement introduce
$\tilde{Q}(t,w)=\Bigl{\{}(z^{0},z)\in\mathbb{R}\times\mathbb{R}^{K(d+1)+1}\Bigm{|}\;$ $z^{0}\geq\ell(w)+\psi(w,u),$
$z=F(w,u),\;$ $u\in U(t,w)\Bigr{\}}.$ (8)
Because $u\mapsto(\ell(w)+\psi(w,u),F(w,u))$ is affine and $U(t,w)$ is convex, each $\tilde{Q}(t,w)$ is convex in $(z^{0},z)$.
All the hypotheses of [27, Th. 1] are therefore verified: the state set is compact and forward invariant, $F$ is Carathéodory, the admissible control map is compact, convex and upper semi-continuous, the integrand is continuous and convex in $u$, and the sets $\tilde{Q}(t,w)$ are convex. Consequently, an optimal admissible pair $(w^{\ast},u^{\ast})$ exists for problem (7). ∎
III Necessary Conditions for Optimality
@@
$H$ $\coloneqq-\sum_{i=1}^{K}\psi^{s}_{i}\left(s^{(i)}\left(f_{i}(\boldsymbol{x})+u_{i}\right)\right)$ (9)
$\quad+\sum_{i=1}^{K}\langle\psi^{x^{(i)}},s^{(i)}f_{i}(\boldsymbol{x})e_{1}+g(x^{(i)})\rangle+\psi^{V}\sum_{i=1}^{K}u_{i}n_{i}s^{(i)}$
@@
$\dot{\psi}^{x^{(i)}}$ $=\sum_{j=1}^{K}(\psi_{j}^{s}-\psi_{1}^{x^{(j)}})s^{(j)}\left(\frac{\partial f_{j}(\boldsymbol{x})}{\partial x^{(i)}}\right)^{\top}$
$\quad-\psi^{x^{(i)}}D_{x}g(x^{(i)})+\lambda_{0}n_{i}\nabla f_{0}\left(x^{(i)}\right)^{\top},$
@@ IV Main Results
Intuitively, $\lvert\psi^{s}_{i}\rvert<\psi^{x^{(i)}}_{1}$ since, under the shadow-price interpretation [12], infection incurs a higher cost than susceptibility. Lemma 1 proves this fact rigorously under the following simplifying assumptions.
Assumption 1
@@
    The state-dependent cost is nontrivial. In particular, if we define
    $\begin{split}I&\coloneqq\Bigl{\{}j\in\{2,\dots,d\}\;\Bigm{|}\;\exists\,m\geq 1,(i_{k})_{k=0,\dots,m}\text{ s. t. }\\ &\quad\quad i_{0}=i,i_{m}=j,\text{ and }G_{i_{k+1},i_{k}}>0,\text{ for all }k\Bigr{\}}.\end{split}$
@@
Proof:
See the Appendix (page Proofs of Auxiliary Results). ∎
@@
Theorem 2 (Optimal vaccination structure)
@@
V Numerical experiments
We validate our theoretical findings by simulating optimal vaccination strategies in three cases (networks with $3$, $5$, and $8$ cities, each city following a simple SIR model with daily commuting between them). All simulations share a 28-day horizon ($T=28$), recovery rate $\gamma=1/7$ day-1, and home-stay fraction $\alpha=0.64$ (workers spend $1-\alpha$ of each week in the destination city). City-specific transmission rates $\beta_{i}\in[0.10,0.35]$, populations, shipment limits $V^{\max}_{w}$, and vaccination capacities $v_{i}^{\max}$. The specific values are given in the Appendix (page Simulation Parameters).
@@ V Numerical experiments
We further test the impact of nonzero and linear migration terms $q_{j}$ (violating Assumption 1) in two experiments with $K=3$ and $K=5$ cities. Figure 2 indicates that the adjoint variables are strictly increasing, which corroborates Lemma 1. This result suggests that even with nonzero $q_{j}$, the bang–bang structure remains intact, and Theorem 2’s result can also be observed. We also solved the problem, but with a quadratic vaccination cost. Figure 3 indicates that the quadratic objective yields a smoother control that approximates the bang-bang profile; the resulting cost difference is negligible (bottom-right panel). The control resulting from the quadratic cost is harder to implement in practice than the bang-bang given in the linear-cost case.
Figure 2: Optimal vaccination control profiles and adjoint variables $\psi^{s}$ for two network structures with $K=3$, and $5$ in the presence of linear and non-zero functions $q_{j}$.
@@ VI Conclusion
Now we present the proof to Lemma 1.
Proof:
Define the variables $z^{(i)}(t)\in\mathbb{R}$, $y^{(i)}(t)\in\mathbb{R}^{d}$, $w^{(i)}(t)\in\mathbb{R}^{d}$, and $\varphi^{(i)}(t)\in\mathbb{R}^{d}$ in the following way: $z^{(i)}(t)\coloneqq\psi_{i}^{s}(T-t)-\psi_{1}^{x^{(i)}}(T-t)$, $y^{(i)}(t)\coloneqq-\psi^{x^{(i)}}(T-t)^{\top}$, $w^{(i)}(t)\coloneqq\sum_{j=1}^{K}z^{(j)}(t)s^{(j)}(T-t)\beta_{ij}+cn_{i},$ $\varphi^{(i)}(t)\coloneqq w^{(i)}(t)+G^{T}y^{(i)}(t)$, and note that the initial conditions are $z^{(i)}(0)=y^{(i)}(0)=0$, $w^{(i)}(0)=cn_{i}$ and $\varphi^{(i)}(0)=cn_{i}$. Using the definitions, these variables satisfy the system
@@
Simulation Parameters
All experiments share the following global parameters:
  • Time horizon: $T=28$ days
  • Recovery rate: $\gamma=1/7$ day-1
  • Home-stay fraction: $\alpha=0.64$
  • Weekly vaccine shipment: $V_{w}^{\max}=\{1,2,3,4\}/30$
  • Cost coefficients: $c_{v}=0.01$, $c_{h}=100$
  • Solver tolerances: $\texttt{RTOL}=\texttt{OTOL}=10^{-6}$
The following parameters vary by experiment:
3 Cities.
[…]
  • Normalized populations: [0.83, 0.083, 0.083]
[…]
  • $v_{i}^{\max}=0.3/T$
  • Commuting matrix:
    $\begin{bmatrix}0.9&0.05&0.05\\ 0.45&0.45&0.10\\ 0.45&0.10&0.45\end{bmatrix}$
5 Cities.
[…]
  • Normalized populations: [0.5, 0.3, 0.1, 0.05, 0.05]
  • $s_{0}=[0.97,\ldots,0.97]$
  • $i_{0}=[0.01,\ldots,0.01]$
[…]
  • Commuting matrix:
    $\begin{bmatrix}0.8&0.05&0.05&0.05&0.05\\ 0.1&0.7&0.1&0.05&0.05\\ 0.05&0.1&0.7&0.1&0.05\\ 0.05&0.05&0.1&0.7&0.1\\ 0.05&0.05&0.05&0.1&0.75\end{bmatrix}$
8 Cities.
  • $\beta=[0.35,0.3,0.25,0.2,0.15,0.1,0.05,0.04]$
  • Normalized populations: [0.46, 0.26, 0.13, 0.06, 0.03, 0.02, 0.007, 0.007]
  • $s_{0}=[0.98,\ldots,0.98]$
  • $i_{0}=[0.005,\ldots,0.005]$
  • $v_{i}^{\max}\!=\![0.5,0.48,0.46,0.44,0.42,0.4,0.38,0.36]/T$
  • Commuting matrix:
    $\begin{bmatrix}0.7&0.1&0.05&0.05&0.03&0.03&0.02&0.02\\ 0.1&0.7&0.1&0.02&0.02&0.02&0.02&0.02\\ 0.05&0.1&0.75&0.02&0.02&0.02&0.02&0.02\\ 0.05&0.02&0.02&0.75&0.05&0.05&0.03&0.03\\ 0.03&0.02&0.02&0.05&0.8&0.03&0.02&0.03\\ 0.03&0.02&0.02&0.05&0.03&0.8&0.03&0.02\\ 0.02&0.02&0.02&0.03&0.02&0.03&0.85&0.01\\ 0.02&0.02&0.02&0.03&0.03&0.02&0.01&0.85\end{bmatrix}$